OpenStudy (anonymous):

A ball is dropped from a height h above the ground . It loses 50% of its velocity on every impact with the ground before bouncing up . What is the total distance travelled by the ball when it finally comes to rest on the gound?

OpenStudy (experimentx):

use the property of geometric series

OpenStudy (anonymous):

Though, you have to make the assumption that eventually reducing the velocity by half each time will yield a velocity close enough to zero as to be negligible. ;-)

OpenStudy (yash2651995):

yes that's what we do.. geometric progression h+h/2+h/2+(h/2)/2+(h/2)/2+(h/4)/2+(h/4)/2...... till infinity.. now maths is only thing left.. try to visualize it.. where ^this thing came from,

OpenStudy (anonymous):

I think this wuestion is beyond my level

OpenStudy (anonymous):

i have nt studied geometric progression!

OpenStudy (anonymous):

Try choosing a value for h (like 100m) and examining a couple bounces using basic kinematics to see the pattern.

OpenStudy (anonymous):

any other way

OpenStudy (anonymous):

Use the first fall to find the final velocity on impact and take half of that to use as the upward velocity for the first bounce - find the new h. Do it again. Look for the pattern. Eventually, the new h will get so small that you'll reasonably call it zero, so you stop.

OpenStudy (experimentx):

\[ d = h \frac{( 1-(1/2)^n)}{1-1/2} = \frac{h}{1-1/2} \text { this is the easiest possible solution}\] when n tends to infinity, (1/2)^n tends to zero ..

OpenStudy (anonymous):

Only thing bugging me about this is that it said the velocity is reduced by half not the height.

OpenStudy (anonymous):

i.e. half the velocity means one-forth the kinetic energy which means one-forth the potential energy which means one-forth of the height each time..

OpenStudy (yash2651995):

ya .. ya.. sorry.. v= sqrt(2gh) v'=v/2=sqrt(2gh')=(sqrt(2gh))/2=>h'=h/4

OpenStudy (anonymous):

The formula for each height as CliffSedge noticed is h(n) = h0(1/2)^(2n) {1} I checked that using kinematics. The ball was released from initial height of h0 so the distance is d = h0 + something every time the ball passes through 2 h(n) of path before it rebounds from the floor. So "something" is \[\sum_{n=1}^{infinity} (2*h(n))\] using {1} we get \[d = h_{0} + \sum_{n=1}^{infinity} 2*h_{0} (1/2)^{2n} \] or \[d = h_{0} + 2h_{0}\sum_{n=1}^{infinity} (1/2)^{2n} \] using wolfram we calculate the thing http://www.wolframalpha.com/input/?i=sum+of&a=*C.sum+of-_*Calculator.dflt-&f2=+%281%2F2%29%5E%282*x%29&x=4&y=10&f=Sum.sumfunction_+%281%2F2%29%5E%282*x%29&f3=1&f=Sum.sumlowerlimit_1&f4=infinity&f=Sum.sumupperlimit_infinity&a=*FVarOpt.1-_**-.***Sum.sumvariable---.*- \[d = h_{0}(1+2/3)=h_{0}*5/3\] Is the answer correct? @openstudy1